IF the national income accounting equation is Y=C+I+G+E-M

Where: C=8+0.6Yd
I=25
G=10
E=70
M=15+02Y
T=7+0.3Y
R=25-0.3Y

1.1 Determine the economy's national income/output
1.2 Determine the economy's multiplier applicable to government spending and interpret its meaning
1.3 Use the multiplier applicable to exports and explain how a N$100 billion decline in the demand for exports could affect the following variables:

i) GDP
ii) Balance of trade
iii) Government budget
1
Expert's answer
2015-03-25T10:00:38-0400
(a) As national income (or GDP) Y = C + I + G + (E - M), then: Y = 8 + 0.6Y + 25 + 10 + 70 - 15 - 0.2Y 0.6Y = 98 Y = 163.33
(b) The fiscal multiplier (m) is the ratio of national income to government spending that causes it. In our case m = G/Y = 10/163.33 = 0.06. This ratio shows how much the increase in government spending may cause the national income to increase.
(c) The foreign trade multiplier also known as export multiplier operates like the investment multiplier of Keynes. It may be defined as the amount by which national income of a nation will be raised by a unit increase in domestic investment on exports. As exports increase there is an increase in the income of all persons associated with the exports industries. In our case m = E/Y = 70/163.33 = 0.43 So, a $100 billion decline in the demand for exports will decrease the total GDP and balance of trade, and also will decrease government budget, as less taxes will be got from exports